# Neural Graph Generator: Feature-Conditioned Graph Generation using Latent Diffusion Models

### v1

@@ Appendix A Dataset Details
Barabási–Albert random graph
250,136 25.01%
Watts–Strogatz small-world graph
204,280 20.43%
stochastic block model graph
125,468 12.55%
Erdős-Rényi graph
122,568 12.26%
dual Barabási–Albert random graph
122,568 12.26%
extended Barabási–Albert model graph
122,568 12.26%
Newman-Watts-Strogatz small-world graph
122,567 12.26%
Holme and Kim algorithm for growing graphs with powerlaw degree distribution and approximate average clustering
122,567 12.26%
Random Lobster graph
81,713 8.17%
random partition graph
81,713 8.17%
Random $d$-degree regular graph
7,000 0.7%
Lollipop graph
4,145 0.41%
Path graph
91 0.01%
Cycle graph
91 0.01%
Star graph
91 0.01%
Wheel graph
91 0.01%
Ladder graph
46 0.0046%
@@
# nodes
[2,100]
Count of individual vertices or entities in a graph.
# edges
[1,4950]
Count of connections or relationships between nodes in the graph.
Density
[0.01,1]
A measure of how interconnected a graph is, calculated as the number of actual edges over the total possible edges.
Min. degree
[1,99]
Minimum degree is the smallest degree among all nodes.
Max. degree
[1,99]
Maximum degree is the largest degree among all nodes.
Avg. degree
[1,99]
Mean degree across all nodes.
Assortativity coefficient
[-1,1]
A measure of the tendency of nodes to connect with others of similar degree. Positive values suggest a preference for connections between nodes of similar degrees.
# triangles
[0, 161700]
Number of sets of three nodes that form a closed loop (triangle) in the network.
Avg. # triangles formed by an edge
[0, 98]
Average number of triangles each edge participates in.
Max. # triangles formed by an edge
[0, 4851]
Highest number of triangles a single edge is part of.
Avg. local clustering coefficient
[0, 1]
A measure of the extent to which nodes in a neighborhood tend to form clusters or cliques.
Global clustering coefficient
[0, 1]
A measure of the overall tendency of nodes to form clusters in the entire graph, reflecting the global structure of the network.
Max. $k$-core
[0, 99]
Largest subgraph where each node has at least $k$ neighbors within the subgraph.
# communities
[0, 50]
Number of clusters of nodes (communities) in the graph.
Diameter
[0, 99]
Maximum distance between any pair of nodes in the graph.

### v2

@@ Appendix A Dataset Details
Barabási–Albert random graph 250,136 25.01%
Watts–Strogatz small-world graph 204,280 20.43%
stochastic block model graph 125,468 12.55%
Erdős-Rényi graph 122,568 12.26%
dual Barabási–Albert random graph 122,568 12.26%
extended Barabási–Albert model graph 122,568 12.26%
Newman-Watts-Strogatz small-world graph 122,567 12.26%
Holme and Kim algorithm for growing graphs with powerlaw degree distribution and approximate average clustering 122,567 12.26%
Random Lobster graph 81,713 8.17%
random partition graph 81,713 8.17%
Random $d$-degree regular graph 7,000 0.7%
Lollipop graph 4,145 0.41%
Path graph 91 0.01%
Cycle graph 91 0.01%
Star graph 91 0.01%
Wheel graph 91 0.01%
Ladder graph 46 0.0046%
@@
# nodes [2,100] Count of individual vertices or entities in a graph.
# edges [1,4950] Count of connections or relationships between nodes in the graph.
Density [0.01,1] A measure of how interconnected a graph is, calculated as the number of actual edges over the total possible edges.
Min. degree [1,99] Minimum degree is the smallest degree among all nodes.
Max. degree [1,99] Maximum degree is the largest degree among all nodes.
Avg. degree [1,99] Mean degree across all nodes.
Assortativity coefficient [-1,1] A measure of the tendency of nodes to connect with others of similar degree. Positive values suggest a preference for connections between nodes of similar degrees.
# triangles [0, 161700] Number of sets of three nodes that form a closed loop (triangle) in the network.
Avg. # triangles formed by an edge [0, 98] Average number of triangles each edge participates in.
Max. # triangles formed by an edge [0, 4851] Highest number of triangles a single edge is part of.
Avg. local clustering coefficient [0, 1] A measure of the extent to which nodes in a neighborhood tend to form clusters or cliques.
Global clustering coefficient [0, 1] A measure of the overall tendency of nodes to form clusters in the entire graph, reflecting the global structure of the network.
Max. $k$-core [0, 99] Largest subgraph where each node has at least $k$ neighbors within the subgraph.
# communities [0, 50] Number of clusters of nodes (communities) in the graph.
Diameter [0, 99] Maximum distance between any pair of nodes in the graph.